OpenStudy (anonymous):

solve the system: x+4=0 y=x+1

OpenStudy (kuoministers):

so we need to find x and y?

OpenStudy (kuoministers):

first we sub x into y=x+1 y=-4 +1 y=-3

OpenStudy (kuoministers):

x = -4

OpenStudy (kuoministers):

Give a best answer if you think any answer helped you!! :D

OpenStudy (kuoministers):

mhm

OpenStudy (kuoministers):

thats great...